# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 5.3.1 Healthcare domain-enhanced LLMs
Likewise, Yang et al. (2023c) introduce Zhongjing, the first Chinese medical language model based on LLaMA, which utilizes a comprehensive training pipeline and a multi-turn medical dialogue dataset. Specifically, it is enhanced with a multi-turn medical dialogue dataset called CMtMedQA, consisting of 70,000 authentic doctor-patient dialogues, enabling complex dialogue and proactive inquiry. The backbone model used is Ziya-LLaMA-13B-v1, and the evaluation dataset is CMtMedQA and huatuo-26M (Li et al., 2023b).
@@
Similarly, Guan et al. (2023) introduce CohortGPT, a model that uses LLMs for participant recruitment in clinical research by classifying complex medical text into disease labels. CohortGPT enhances ChatGPT performance with the use of a knowledge graph as auxiliary information and a CoT sample selection strategy. The tasks involve IU-RR (Preparing a collection of radiology examinations for distribution and retrieval) and MIMIC-CXR (Mimic-cxr, a de-identified publicly available database of chest radiographs with free-text reports). Li et al. (2023f) introduce ChatDoctor, a refined LLaMA-based model. It is fine-tuned using a dataset of 100,000 patient-doctor dialogues, and equipped with a self-directed information retrieval mechanism.
@@
Jin et al. (2023) present GeneGPT, a method for teaching LLMs to answer genomics questions using NCBI Web APIs. Zakka et al. (2023) introduce Almanac, a LLM with retrieval capabilities for medical guideline recommendations. Lastly, Liu et al. (2023d) introduce MolXPT, a unified language model adept in molecular property prediction and molecular generation.
5.3.2 Legal domain enhanced LLMs
@@
Huang et al. (2023a) address the challenge of enhancing LLMs like LLaMA for domain-specific tasks, particularly in the legal domain, by injecting domain knowledge during continual training, designing appropriate supervised finetune tasks and incorporating a retrieval module to improve factuality during text generation. They release their data and model for further research in Chinese legal tasks.
@@
5.3.3 Finance Domain-enhanced LLMs
@@
Li et al. (2023e) introduce EcomGPT, a language model tailored for E-commerce scenarios, trained on the newly created EcomInstruct dataset, which consists of 2.5 million instruction data spanning various E-commerce tasks and data types. The dataset covers product information, user reviews, and more. It defines atomic tasks and Chain-of-Task tasks to enable comprehensive training for E-commerce scenarios. The backbone model used is BLOOMZ, which is fine-tuned with the EcomInstruct dataset. The evaluation dataset includes 12 tasks, encompassing classification, generation, extraction, and other E-commerce-related tasks.
@@
5.3.4 Other Domain-Enhanced LLMs
@@
6 Conclusion
Throughout this survey, we have systematically explored the intricate landscape of factuality issues within large language models (LLMs). We began by defining the concept of factuality (Sec 2.2) and proceeded to discuss its broader implications (Sec 2.3). Our journey took us through the multifaceted realm of factuality evaluation, encompassing benchmarks (Sec 3.2), metrics (Sec 3.1), specific evaluation studies (Sec 3.3), and domain-specific evaluations (Sec 3.4). We then delved deeper, probing the intrinsic mechanisms that underpin factuality in LLMs (Sec 4). Our exploration culminated in the discussion of enhancement techniques, both for pure LLMs (Sec 5.1) and retrieval-augmented LLMs (Sec 5.2), with a special focus on domain-specific LLM enhancements (Sec 5.3).

### v3

@@ 5.4. Healthcare domain-enhanced LLMs
Likewise, Yang et al. (2023c) introduce Zhongjing, the first Chinese medical language model based on LLaMA, which utilizes a comprehensive training pipeline and a multi-turn medical dialogue dataset. Specifically, it is enhanced with a multi-turn medical dialogue dataset called CMtMedQA, consisting of 70,000 authentic doctor-patient dialogues, enabling complex dialogue and proactive inquiry. The backbone model used is Ziya-LLaMA-13B-v1, and the evaluation dataset is CMtMedQA and huatuo-26M (Li et al., 2023f).
@@
Similarly, Guan et al. (2023) introduce CohortGPT, a model that uses LLMs for participant recruitment in clinical research by classifying complex medical text into disease labels. CohortGPT enhances ChatGPT performance with the use of a knowledge graph as auxiliary information and a CoT sample selection strategy. The tasks involve IU-RR (Preparing a collection of radiology examinations for distribution and retrieval) and MIMIC-CXR (Mimic-cxr, a de-identified publicly available database of chest radiographs with free-text reports). Li et al. (2023b) introduce ChatDoctor, a refined LLaMA-based model. It is fine-tuned using a dataset of 100,000 patient-doctor dialogues, and equipped with a self-directed information retrieval mechanism.
@@
Jin et al. (2023) present GeneGPT, a method for teaching LLMs to answer genomics questions using NCBI Web APIs. Zakka et al. (2023) introduce Almanac, a LLM with retrieval capabilities for medical guideline recommendations. Lastly, Liu et al. (2023f) introduce MolXPT, a unified language model adept in molecular property prediction and molecular generation.
5.5. Legal domain enhanced LLMs
@@
Huang et al. (2023c) address the challenge of enhancing LLMs like LLaMA for domain-specific tasks, particularly in the legal domain, by injecting domain knowledge during continual training, designing appropriate supervised finetune tasks and incorporating a retrieval module to improve factuality during text generation. They release their data and model for further research in Chinese legal tasks.
@@
5.6. Finance Domain-enhanced LLMs
@@
Li et al. (2023c) introduce EcomGPT, a language model tailored for E-commerce scenarios, trained on the newly created EcomInstruct dataset, which consists of 2.5 million instruction data spanning various E-commerce tasks and data types. The dataset covers product information, user reviews, and more. It defines atomic tasks and Chain-of-Task tasks to enable comprehensive training for E-commerce scenarios. The backbone model used is BLOOMZ, which is fine-tuned with the EcomInstruct dataset. The evaluation dataset includes 12 tasks, encompassing classification, generation, extraction, and other E-commerce-related tasks.
@@
5.7. Other Domain-Enhanced LLMs
@@
Bi et al. (2023) present OceanGPT, the first-ever LLM in the ocean domain, which is expert in various ocean science tasks. They also propose a novel framework called DoInstruct to automatically obtain a large volume of ocean domain data. OceanGPT is evaluated in OceanBench and shows a higher level of knowledge expertise for oceans science tasks.
@@
6. Conclusion
Throughout this survey, we have systematically explored the intricate landscape of factuality issues within large language models (LLMs). We began by defining the concept of factuality (Sec 2.2) and proceeded to discuss its broader implications (Sec 2.3). Our journey took us through the multifaceted realm of factuality evaluation, encompassing benchmarks (Sec 3.2), metrics (Sec 3.1), specific evaluation studies (Sec 3.3), and domain-specific evaluations (Sec 3.4). We then delved deeper, probing the intrinsic mechanisms that underpin factuality in LLMs (Sec 4). Our exploration culminated in the discussion of enhancement techniques, both for standalone LLMs (Sec 5.1) and retrieval-augmented LLMs (Sec 5.2), with a special focus on domain-specific LLM enhancements (Sec 5.3).